# HiLoMix: Robust High- and Low-Frequency Graph Learning Framework for Mixing Address Association

### v1

@@ Abstract
As mixing services are increasingly being exploited by malicious actors for illicit transactions, mixing address association has emerged as a critical research task. A range of approaches have been explored, with graph-based models standing out for their ability to capture structural patterns in transaction networks. However, these approaches face two main challenges: label noise and label scarcity, leading to suboptimal performance and limited generalization. To address these, we propose HiLoMix, a graph-based learning framework specifically designed for mixing address association. First, we construct the Heterogeneous Attributed Mixing Interaction Graph (HAMIG) to enrich the topological structure. Second, we introduce frequency-aware graph contrastive learning that captures complementary structural signals from high- and low-frequency graph views. Third, we employ weak supervised learning that assigns confidence-based weighting to noisy labels. Then, we jointly train high-pass and low-pass GNNs using both unsupervised contrastive signals and confidence-based supervision to learn robust node representations. Finally, we adopt a stacking framework to fuse predictions from multiple heterogeneous models, further improving generalization and robustness. Experimental results demonstrate that HiLoMix outperforms existing methods in mixing address association.
@@ Introduction
Machine learning-based approaches automate the association of mixing addresses by training predictive models that learn implicit transaction patterns from blockchain data. (hu2023bert4eth) employed a pre-trained Transformer (vaswani2017attention) to encode transaction sequences of Ethereum accounts, capturing temporal and contextual behavior patterns. Subsequently, (hu2024zipzap) achieved both parameter and computational efficiency using frequency-aware compression and asymmetric training compared (hu2023bert4eth). However, their approaches overlook the structural relationships among accounts, which are crucial for mixing address association. (du2023breaking) modeled address interactions as a Mixing Interaction Graph (MIG) and formulated the task as a node-pair link prediction problem. By applying GraphSAGE (hamilton2017inductive), their approach captures the topological structures and interaction dependencies among accounts. Furthermore, (beres2021blockchain) applied classic graph representation learning for Ethereum address association. (zhong2024bitlink) developed a deep neural network to link disjoint Bitcoin address clusters in a self-supervised manner. Although these approaches are not specifically designed for mixing services, they can be readily adapted to our task due to the similarity in task formulation. However, machine learning models rely on labeled data, which are unavailable for mixing address associations. As a workaround, heuristic rules are employed to generate address associations but inevitably introduce noise, compromising both training and evaluation. To the best of our knowledge, the impact of such label noise has not been formally investigated in this context.
@@
As depicted in Figure 2, HiLoMix comprises four key components: (1) HAMIG construction, which encodes diverse account interactions into a heterogeneous graph to enrich the topological structure; (2) Frequency-aware graph contrast learning, which contrasts node representations derived from high- and low-frequency views to capture robust structural patterns; (3) Confidence-based supervision, which assigns adaptive weights to noisy labels based on estimated reliability to mitigate supervision noise; (4) Mixing address association, which employs a stacking framework to integrate the knowledge of multiple models for final prediction.
@@
(du2023breaking) constructed the Mixing Interaction Graph (MIG) to capture the topological structure of account associations, but ignored the issue of structural sparsity resulting from limited association links. To address the challenge of ineffective message propagation caused by this sparsity, we augment MIG by incorporating mixing transaction interactions, resulting in a heterogeneous account interaction graph that significantly enriches the graph structure. We define this enhanced structure as the Heterogeneous Attributed Mixing Interaction Graph (HAMIG), which captures both the topological structure of account interactions and attribute information of account nodes. Unlike MIG, HAMIG is formulated as an undirected heterogeneous graph $\mathcal{G}=(\mathcal{V}_{a},\mathcal{V}_{t},\mathcal{E}_{at},\mathcal{E}_{aa},\mathcal{X})$, where $\mathcal{V}_{a}$ is the set of account nodes interacting with Tornado Cash and $\mathcal{V}_{t}$ is the set of smart contract nodes within Tornado Cash, $\mathcal{E}_{at}=\{(v_{i},v_{j})|v_{i}\in\mathcal{V}_{a},v_{j}\in\mathcal{V}_{t}\}$ represents the set of mixing transaction edges, $\mathcal{E}_{aa}=\{(v_{i},v_{j})|v_{i},v_{j}\in\mathcal{V}_{a}\}$ denotes account association edges, and $\mathcal{X}$ is the node feature matrix constructed from raw mixing transaction data. Details of HAMIG construction can be found in Appendix.
@@
where $\parallel$ denotes vector concatenation, and the sigmoid function ensures that $s_{ij}$ is bounded within $[0,1]$. Based on estimated edge smoothness scores, HAMIG $\mathcal{G}$ is decomposed into two frequency-specific graph views: a low-frequency view $\mathcal{G}^{(LF)}$ and a high-frequency view $\mathcal{G}^{(HF)}$. Both views share the same node set, while the adjacency matrix $A$ is partitioned into $A^{(LF)}$ and $A^{(HF)}$ as follows:
@@
$\mathcal{H}_{l}^{(LF)}=\left(I+\tilde{A}^{(LF)}\right)\mathcal{H}_{l-1}^{(LF)}$ (3)
@@
$\mathcal{H}_{l}^{(HF)}=\left(I-\alpha\cdot\tilde{A}^{(HF)}\right)\mathcal{H}_{l-1}^{(HF)}$ (4)
Here, $\tilde{A}^{(LF)}$ and $\tilde{A}^{(HF)}$ denote the normalized adjacency matrices for the low- and high-frequency views, respectively; $l\in\{1,...,L\}$ is the layer index; $\alpha$ controls the high-pass filter intensity; and $\mathcal{H}_{L}^{(LF)}$ and $\mathcal{H}_{L}^{(HF)}$ (i.e., $\mathcal{H}^{(LF)}$ and $\mathcal{H}^{(HF)}$) are the final node embeddings from the two views.
By applying low-pass and high-pass GNNs to the low- and high-frequency graph views, respectively, we obtain node representations that emphasize distinct frequency semantics. These representations, $\mathcal{H}^{(LF)}$ and $\mathcal{H}^{(HF)}$, are then contrasted using InfoNCE loss (oord2018representation) to facilitate frequency-aware representation learning.
$\mathcal{L}^{(LF)}\left(h_{i}^{(LF)},h_{i}^{(HF)}\right)=-\log\frac{s\left(h_{i}^{(LF)},h_{i}^{(HF)}\right)}{\sum_{i\neq j}s\left(h_{i}^{(LF)},h_{j}^{(HF)}\right)}$ (5)
$\mathcal{L}^{(HF)}\left(h_{i}^{(HF)},h_{i}^{(LF)}\right)=-\log\frac{s\left(h_{i}^{(HF)},h_{i}^{(LF)}\right)}{\sum_{i\neq j}s\left(h_{i}^{(HF)},h_{j}^{(LF)}\right)}$ (6)
@@
$\mathcal{L}_{mul}^{ij}$ $=-y_{ij}\left[\log\left(p_{ij}^{(LF)}\right)+\log\left(p_{ij}^{(HF)}\right)\right]$
$=-y_{ij}\log\left(p_{ij}^{(LF)}\cdot p_{ij}^{(HF)}\right)$ (8)
The mutual loss gauges predictive confidence: $\mathcal{L}_{mul}^{ij}$ has a low value when both networks confidently and correctly predict the presence of edge $e_{ij}$.
@@
Deep neural networks exhibit a memorization effect, where clean and simple patterns are learned in the early stages of training, while noisy labels are gradually overfitted in later epochs (arpit2017closer). To adaptively manage label noise during training, we adopt the small-loss criterion (han2018co), which selects clean samples based on training losses. At epoch $t$, we define a threshold to distinguish clean from noisy labels as follows:
$\mathcal{L}_{thre}^{t}=\text{Percentile}\left(\mathcal{L}_{mul}^{ij},100-\frac{50\cdot t}{T_{\text{max}}}\right)$ (9)
where $T_{\text{max}}$ denotes the total number of training epochs, $0.5\leq\left(100-\frac{50\cdot t}{T_{\text{max}}}\right)<1$ ensures that at least half of the labels are retained as clean, $\text{Percentile}\left(\mathcal{L},p\right)$ returns the value below which pp% of the loss values in $\mathcal{L}$ fall. This dynamic threshold allows more samples to be included during the early stages for sufficient training and gradually reduces the number of selected instances as training progresses.
We define $\mathcal{L}_{thre}^{avg}=\text{Average}(\mathcal{L}_{mul}^{ij})$ to further ensure that small loss samples are clean regardless of their relative rank. Then clean and noisy label sets can be roughly divided as follows:
@@
where the confidence score $\mu(e_{ij})$ is defined as $\mu(e_{ij})=\sqrt{p_{ij,c}^{(LF)}\cdot p_{ij,c}^{(HF)}}$. A higher value of $\mu(e_{ij})$ suggests a greater likelihood that $e_{ij}$ is mislabeled, yet correctly predicted. The threshold $Thre$ is designed to be dynamic: stricter in the early stages of training and gradually relaxed as training progresses.
Given the previously defined sets $\mathcal{E}_{cl}$ and $\mathcal{E}_{cf}$, we denote the remaining label set as $\mathcal{E}_{re}$, from which no useful information can be reliably extracted.
@@
$\mathcal{L}_{sup}=\frac{1}{|\mathcal{E}_{aa}|}\sum_{e_{ij}\in\mathcal{E}_{aa}}\xi(e_{ij})\,\hat{y}_{ij}\log\left(p_{ij}^{(LF)}\cdot p_{ij}^{(HF)}\right)$ (14)
@@
where $\xi(\cdot)$ serves as a confidence score, assigning each label a different weight to guide the model toward learning from more reliable supervision. Finally, the total loss used to update both the high-pass and low-pass GNNs is formulated as follows:
@@
We collect 554,589 raw Tornado Cash mixing transactions spanning from 2019-12-16 to 2025-01-17. After data cleaning and preprocessing, 371,181 valid transactions are retained. Following the heuristic rules proposed in (wu2022tutela; wang2023zero), we identify 4,074 unique address associations. By modeling these associations and transaction interactions as edges, we construct a heterogeneous HAMIG comprising 106,982 account nodes and 375,255 edges of two types. Further details of dataset construction are provided in the Appendix.
@@
To comprehensively evaluate the performance of HiLoMix, we adopt four metrics across two categories: classification and ranking. Specifically, F1-score and AUC are used to evaluate classification performance, while MRR and Hits@K assess ranking effectiveness. Details of these metrics are provided in the Appendix.
@@
We compare HiLoMix with 11 baseline models, grouped into three categories: (1) Graph representation learning methods: DeepWalk (perozzi2014deepwalk), Node2Vec (grover2016node2vec), GCN (kipf2016semi), GAT (velickovic2017graph), GIN (xu2018powerful), and GraphSAGE (hamilton2017inductive). Since the mixing address association is formulated as a node-pair link prediction task in HAMIG, classical graph representation learning models serve as strong baselines. (2) Robust learning methods: Co-teaching (han2018co), Co-teaching+ (yu2019does), and NRGL (wu2024robust). Given the central limitation of label noise in our task, we include both classical and state-of-the-art robust learning approaches for comparison. (3) Mixing address association methods: BERT4ETH (hu2023bert4eth) and MixBroker (du2023breaking). These methods are designed specifically for the mixing address association task.
@@
Mixing address association under the classification setting. HiLoMix achieves an AUC of 0.9137 and a $\text{F}_{1}$ score of 0.8382, outperforming all baselines with relative improvements ranging from 7.34% to 29.09% in AUC and 5.69% to 22.47% in $\text{F}_{1}$, demonstrating its effectiveness in the classification setting. Classic graph representation learning methods perform the worst, primarily due to their vulnerability to label noise and graph sparsity resulting from scarce associations. NRGL achieves the second highest AUC, validating the effectiveness of explicit label noise handling. MixBroker attains the second-best $\text{F}_{1}$ score, benefiting from task-specific graph modeling that captures user interaction structures. In contrast, BERT4ETH underperforms standard GNNs, highlighting the importance of incorporating structural information.
Mixing address association under the ranking setting. HiLoMix outperforms the second-best model by 15.61%, 8.13%, 11.18%, and 10.65% in MRR, Hits@3, Hits@5 and Hits@10, respectively. These results indicate that HiLoMix effectively ranks true address associations near the top among all candidates, demonstrating a strong generalization to unseen links. In line with the classification results, the classic graph representation learning methods perform the worst. NRGL ranks second across all ranking metrics, further validating the effectiveness of robust learning strategies for mixing address association task. BERT4ETH leverages sequence modeling to generate expressive account representations; however, its disregard for structural dependencies leads to suboptimal performance. MixBroker, while incorporating user interaction structures through graph modeling, suffers from graph sparsity and structural perturbations.
@@
To systematically validate the effectiveness of each component in HiLoMix, we conduct ablation studies by removing or modifying specific modules. Specifically, “w/o hetero graph” removes all mixing transaction edges, retaining only address association links in HAMIG. “w/o HiLo learning” replaces the high- and low-pass GNNs with two standard GCNs. “w/o con loss” removes the contrastive loss from the overall objective. “w/o label division” disables the confidence-based label division mechanism. “w/o stacking” discards the stacking framework and instead averages the predictions of the high- and low-pass GNNs. “w/o hetero models” excludes the additional heterogeneous base models from the stacking ensemble.
@@
As shown in Table 2, removing any component results in a consistent performance drop across both classification and ranking metrics, validating the complementary contributions of all modules. Among all ablation settings, removing the stacking framework or excluding heterogeneous base models leads to the most substantial degradation: $\text{F}_{1}$ drops by 6.65% and 11.80%, while MRR declines by 32.16% and 35.45%, respectively. These results highlight the effectiveness of the stacking strategy in integrating knowledge from diverse models and underscore the importance of base model heterogeneity. The remaining four ablation variants cause moderate decreases, with $\text{F}_{1}$, AUC, and MRR reduced by 1.72% to 3.57%, 2.18% to 2.58%, and 4.18% to 9.22%, respectively, further validating the necessity of each component. These relatively smaller drops also emphasize the robustness conferred by the stacking framework, which effectively aggregates noise-resistant and complementary knowledge from heterogeneous base models.
@@
To assess the effectiveness of our constructed features, we conduct a comparative study across seven feature variants by progressively removing less important dimensions. Additionally, we incorporate node features from MixBroker for further comparison. As shown in Table 3, HiLoMix achieves the best performance across all metrics when using our 197-dimensional account node features, outperforming the second-best features, MixBroker features, by 1.62%, 1.24%, and 4.27% in $\text{F}_{1}$, AUC, and MRR, respectively. Although the 221-dimensional feature set contains more attributes, its performance is inferior to that of the 197-dimensional set due to reduced interpretability. These results suggest that our 197-dimensional features not only enhance model explainability but also reduce HAMIG’s memory footprint without compromising performance.
@@
To evaluate the stability of our model, we perform a sensitivity analysis on key hyper-parameters in HiLoMix. We begin by evaluating model stability with respect to the number of folds used in cross-validation within the stacking ensemble. As shown in Figure 3, model performance remains generally stable across different fold settings, with $\text{F}_{1}$, AUC, and MRR maintaining relatively high values. The best overall performance is observed with 5-fold cross-validation, which is therefore adopted as the default setting in HiLoMix. In contrast, both fewer and greater numbers of folds lead to slight performance degradation, potentially due to overfitting or unstable ensemble integration.
We further investigate the sensitivity of HiLoMix to two additional hyper-parameters: the loss weighting $\lambda$, which balances the supervised and contrastive learning objectives, and the high-pass filter strength $\alpha$, which modulates the emphasis on high-frequency components. As illustrated in Figure 4, HiLoMix demonstrates robust performance across a wide range of $\lambda$ and $\alpha$ values, with $\text{F}_{1}$ scores ranging from approximately 0.80 to 0.84 and MRR values ranging from 0.42 to 0.46. Both metrics peak around $\lambda=2$ and $\alpha=0.5$, suggesting that a moderate balance between the supervised and contrastive learning objectives and well-calibrated filter strength are critical for optimal performance.
Figure 3: Performance of HiLoMix when training the base models using K-fold cross-validation in the stacking framework.

### v2

@@ Abstract
As mixing services are increasingly being exploited by malicious actors for illicit transactions, mixing address association has emerged as a critical research task. A range of approaches have been explored, with graph-based models standing out for their ability to capture structural patterns in transaction networks. However, these approaches face two main challenges: label noise and label scarcity, leading to suboptimal performance and limited generalization. To address these, we propose HiLoMix, a graph-based learning framework specifically designed for mixing address association. First, we construct the Heterogeneous Attributed Mixing Interaction Graph (HAMIG) to enrich the topological structure. Second, we introduce frequency-aware graph contrastive learning that captures complementary structural signals from high- and low-frequency graph views. Third, we employ weak supervised learning that assigns confidence-based weights to noisy labels. Then, we jointly train high-pass and low-pass GNNs using both unsupervised contrastive signals and confidence-based supervision to learn robust node representations. Finally, we adopt a stacking framework to fuse predictions from multiple heterogeneous models, further improving generalization and robustness. Experimental results demonstrate that HiLoMix outperforms existing methods in mixing address association.
@@ Introduction
Machine learning-based approaches automate the association of mixing addresses by training predictive models that learn implicit transaction patterns from blockchain data. (hu2023bert4eth) employed a pre-trained Transformer (vaswani2017attention) to encode transaction sequences of Ethereum accounts, capturing temporal and contextual behavior patterns. Subsequently, (hu2024zipzap) achieved both parameter and computational efficiency using frequency-aware compression and asymmetric training compared (hu2023bert4eth). However, their approaches overlook the structural relationships among accounts, which are crucial for mixing address association. (du2023breaking) modeled the address interactions as a Mixing Interaction Graph (MIG) and formulated the task as a node-pair link prediction problem. By applying GraphSAGE (hamilton2017inductive), their approach captures the topological structures and interaction dependencies among accounts. Furthermore, (beres2021blockchain) applied classic graph representation learning for address associations. (zhong2024bitlink) developed a deep neural network to link disjoint Bitcoin address clusters in a self-supervised manner. Although these approaches are not specifically designed for mixing services, they can be readily adapted to our task due to the similarity in task formulation. However, machine learning models heavily rely on labeled data, which is unavailable for the mixing address association task. As a workaround, heuristic rules are employed to generate address associations but inevitably introduce noise, compromising both training and evaluation. To the best of our knowledge, the impact of such label noise has not been formally investigated in this context.
@@
As depicted in Figure 2, HiLoMix comprises four key components: (1) HAMIG construction, which encodes diverse account interactions into a heterogeneous graph to enrich the topological structure; (2) Frequency-aware graph contrast learning, which contrasts node representations from high- and low-frequency views to capture robust structural patterns; (3) Confidence-based supervision, which assigns adaptive weights to noisy labels based on estimated reliability to mitigate supervision noise; (4) Mixing address association, which employs a stacking framework to integrate the knowledge of multiple models for final prediction.
@@
(du2023breaking) introduced the Mixing Interaction Graph (MIG) to capture the topological structure of account associations, where each account is modeled as a node and the address associations are represented as edges. Although effective in describing the basic interaction topology, MIG suffers from structural sparsity due to limited address associations, which restricts effective message propagation in graph-based learning. To address this limitation, we extend MIG by incorporating mixing transaction interactions as additional edges, thereby constructing a heterogeneous account interaction graph that significantly enriches the structural connectivity and semantic information of MIG. In addition, we extract a node feature matrix $\mathcal{X}\in\mathbb{R}^{n\times d}$, where $n$ denotes the number of account nodes and $d$ the feature dimension. Each feature vector encodes an account’s interaction behaviors with Tornado Cash mixing contracts, including its interaction frequency, timestamps, and gas price statistics. We define the resulting enhanced structure as the Heterogeneous Attributed Mixing Interaction Graph (HAMIG), which jointly captures both the topological dependencies among accounts and the attribute characteristics of individual nodes. Unlike MIG, HAMIG is formulated as an undirected heterogeneous graph $\mathcal{G}=(\mathcal{V}_{a},\mathcal{V}_{t},\mathcal{E}_{at},\mathcal{E}_{aa},\mathcal{X})$, where $\mathcal{V}_{a}$ denotes the set of account addresses interacting with Tornado Cash, $\mathcal{V}_{t}$ denotes the set of smart contracts within Tornado Cash, $\mathcal{E}_{at}=\{(v_{i},v_{j})\mid v_{i}\in\mathcal{V}_{a},v_{j}\in\mathcal{V}_{t}\}$ represents mixing transaction edges, $\mathcal{E}_{aa}=\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V}_{a}\}$ represents address association edges, and $\mathcal{X}$ is the node feature matrix.
@@
where $h_{i}$ indicates the feature vector of the $i$-th node, $\parallel$ denotes vector concatenation, and the sigmoid function ensures that $s_{ij}$ is bounded within $[0,1]$. Based on estimated edge smoothness scores, HAMIG $\mathcal{G}$ is decomposed into two frequency-specific graph views: a low-frequency view $\mathcal{G}^{(LF)}$ and a high-frequency view $\mathcal{G}^{(HF)}$. Both views share the same node set, while the adjacency matrix $A$ is partitioned into $A^{(LF)}$ and $A^{(HF)}$ as follows:
@@
$\mathcal{H}_{l}^{(LF)}=\left(I+\tilde{A}^{(LF)}\right)\mathcal{H}_{l-1}^{(LF)},\quad l\in\{1,...,L\}$ (3)
@@
$\mathcal{H}_{l}^{(HF)}=\left(I-\alpha\cdot\tilde{A}^{(HF)}\right)\mathcal{H}_{l-1}^{(HF)},\quad l\in\{1,...,L\}$ (4)
Here, $\tilde{A}^{(LF)}$ and $\tilde{A}^{(HF)}$ denote the normalized adjacency matrices for the low- and high-frequency graph views, respectively; $\alpha$ controls the high-pass filter intensity; and $\mathcal{H}_{L}^{(LF)}$ and $\mathcal{H}_{L}^{(HF)}$ (i.e., $\mathcal{H}^{(LF)}$ and $\mathcal{H}^{(HF)}$) are the final node embeddings from the two views.
By applying low-pass and high-pass GNNs to the low- and high-frequency graph views, respectively, we obtain node representations that emphasize distinct frequency semantics. These representations, $\mathcal{H}^{(LF)}$ and $\mathcal{H}^{(HF)}$, are then contrasted using InfoNCE loss (oord2018representation) to facilitate node representation learning.
$\mathcal{L}^{(LF)}\left(h_{i}^{(LF)},h_{i}^{(HF)}\right)=-\log\frac{s\left(h_{i}^{(LF)},h_{i}^{(HF)}\right)}{\sum_{j=1}^{N}s\left(h_{i}^{(LF)},h_{j}^{(HF)}\right)}$ (5)
$\mathcal{L}^{(HF)}\left(h_{i}^{(HF)},h_{i}^{(LF)}\right)=-\log\frac{s\left(h_{i}^{(HF)},h_{i}^{(LF)}\right)}{\sum_{j=1}^{N}s\left(h_{i}^{(HF)},h_{j}^{(LF)}\right)}$ (6)
@@
$\mathcal{L}_{mul}^{ij}$ $=-y_{ij}\left[\log\left(p_{ij,\;y_{ij}}^{(LF)}\right)+\log\left(p_{ij,\;y_{ij}}^{(HF)}\right)\right]$
$=-y_{ij}\log\left(p_{ij,\;y_{ij}}^{(LF)}\cdot p_{ij,\;y_{ij}}^{(HF)}\right)$ (8)
Mutual loss gauges predictive confidence: $\mathcal{L}_{mul}^{ij}$ has a low value when both networks confidently and correctly predict the presence of edge $e_{ij}$.
@@
Deep neural networks exhibit a memorization effect, where clean and simple patterns are learned in the early stages of training, while noisy labels are gradually overfitted in later epochs (arpit2017closer). To adaptively manage label noise during training, we adopt the small-loss criterion (han2018co), which selects clean samples based on training losses. In epoch $t$, we define a threshold to distinguish clean from noisy labels as follows:
$\mathcal{L}_{thre}^{t}=\text{Percentile}\left(\mathcal{L}_{mul}^{ij},1-\frac{t}{2\cdot T_{\text{max}}}\right)$ (9)
where $T_{\text{max}}$ denotes the total number of training epochs, $0.5\leq\left(1-\frac{t}{2\cdot T_{\text{max}}}\right)<1$ ensures that at least half of the labels are retained as clean, $\text{Percentile}\left(\mathcal{L},p\right)$ returns the value below which 100p100*p% of the loss values in $\mathcal{L}$ fall. This dynamic threshold allows more samples to be included during the early stages for sufficient training and gradually reduces the number of selected instances as training progresses.
We define $\mathcal{L}_{thre}^{avg}=\text{Average}(\mathcal{L}_{mul}^{ij})$ to further ensure that small loss samples are clean regardless of their relative rank. Then clean and noisy label sets can be divided as follows:
@@
where the confidence score $\mu(e_{ij})$ is defined as $\mu(e_{ij})=\sqrt{p_{ij,c}^{(LF)}\cdot p_{ij,c}^{(HF)}}$. A higher value of $\mu(e_{ij})$ suggests a greater likelihood that $e_{ij}$ is mislabeled, yet correctly predicted. The threshold $Thre$ is designed to be dynamic: stricter in the early stages of training and gradually relaxed as the training progresses.
Given the previously defined sets $\mathcal{E}_{cl}$ and $\mathcal{E}_{cf}$, we denote the remaining labels as $\mathcal{E}_{re}$, from which no useful information can be reliably extracted.
@@
$\mathcal{L}_{\mathrm{sup}}=-\frac{1}{|\mathcal{E}_{aa}|}\sum_{e_{ij}\in\mathcal{E}_{aa}}\xi(e_{ij})\big(\log p^{(LF)}_{ij,\;\hat{y}_{ij}}+\log p^{(HF)}_{ij,\;\hat{y}_{ij}}\big)$ (14)
@@
where $\xi(\cdot)$ serves as a confidence score, assigning each label a different weight to guide the model towards more reliable supervision. Finally, the total loss used to update both the high-pass and low-pass GNNs is formulated as follows:
@@
We collect 554,589 raw Tornado Cash mixing transactions spanning from 2019-12-16 to 2025-01-17. After data cleaning and preprocessing, 371,181 valid transactions are retained. Following the heuristic rules proposed in (wu2022tutela; wang2023zero), we identify 4,074 unique address associations as labels. By modeling these associations and transaction interactions as edges, we construct HAMIG comprising 106,982 account nodes and 375,255 edges of two types.
Although our dataset is constructed solely from Tornado Cash, we note that this does not compromise generality. Tornado Cash remains the largest and most representative Ethereum-based mixer, accounting for over 95% of the total transaction volume and total value locked (TVL) across all Ethereum mixing services according to (akopovadetecting). Therefore, focusing on Tornado Cash already captures the predominant behavioral and structural patterns of Ethereum-based mixing activities.
@@
To comprehensively evaluate the performance of HiLoMix, we adopt four metrics across two categories: classification and ranking. Specifically, F1-score and AUC are used to evaluate classification performance, while MRR and Hits@K assess ranking effectiveness. The formal definitions of MRR and Hits@K are provided:
$\mathrm{MRR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\mathrm{rank}_{i}}$ (17)
$\mathrm{Hits@K}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\mathrm{rank}_{i}\leq K)$ (18)
where $N$ is the number of positive samples, $\mathrm{rank}_{i}$ denotes the ranking position of the $i$-th positive sample in the candidate list based on predicted probabilities, and $\mathbb{I}(\cdot)$ is the indicator function.
@@
We compare HiLoMix with 11 baseline models, grouped into three categories: (1) Graph representation learning methods: DeepWalk (perozzi2014deepwalk), Node2Vec (grover2016node2vec), GCN (kipf2016semi), GAT (velickovic2017graph), GIN (xu2018powerful), and GraphSAGE (hamilton2017inductive). Since the mixing address association is formulated as a node-pair link prediction task in HAMIG, classical graph representation learning models serve as strong baselines. (2) Robust learning methods: Co-teaching (han2018co), Co-teaching+ (yu2019does) and NRGL (wu2024robust). Given the central limitation of label noise in our task, we include both classical and state-of-the-art robust learning approaches for comparison. (3) Mixing address association methods: BERT4ETH (hu2023bert4eth) and MixBroker (du2023breaking). These methods are designed specifically for the mixing address association task.
Implementation Details:
All implementations are conducted using PyTorch 2.3.1 and DGL 2.4.0. The model is trained for 50 epochs with a batch size of 128, using the Adam optimizer with a learning rate of 0.003. All experiments are performed on a single NVIDIA A100 GPU. For contrastive loss, we use a batch of 128 nodes and set the temperature coefficient at 0.5. For MRR and Hits@K metrics, 50 negative samples are generated for each positive instance. All baselines are implemented following their original papers and fine-tuned to achieve optimal performance.
@@
Mixing address association under the classification setting. HiLoMix achieves an AUC of 0.9137 and a $\text{F}_{1}$ score of 0.8382, outperforming all baselines with relative improvements ranging from 7.34% to 29.09% in AUC and 5.69% to 22.47% in $\text{F}_{1}$, demonstrating its effectiveness in the classification setting. Classic graph representation learning methods perform the worst, primarily due to their vulnerability to label noise and graph sparsity resulting from scarce associations. NRGL achieves the second highest AUC, validating the effectiveness of explicit label noise handling. MixBroker attains the second-best $\text{F}_{1}$ score, benefiting from task-specific graph modeling that captures user interaction structures. In contrast, BERT4ETH underperforms standard GNNs, highlighting the importance of incorporating structural information in address association.
Mixing address association under the ranking setting. HiLoMix outperforms the second-best model by 15.61%, 8.13%, 11.18% and 10.65% in MRR, Hits@3, Hits@5 and Hits@10, respectively. These results indicate that HiLoMix effectively ranks true address associations near the top among all candidates, demonstrating a strong generalization to unseen links. In line with the classification results, the classic graph representation learning methods perform the worst. NRGL ranks second across all ranking metrics, further validating the effectiveness of robust learning strategies for mixing address association task. BERT4ETH leverages sequence modeling to generate expressive account representations; however, its disregard for structural dependencies leads to suboptimal performance. MixBroker, while incorporating user interaction structures through graph modeling, suffers from graph sparsity and structural perturbations.
@@
To systematically validate the effectiveness of each component in HiLoMix, we conduct ablation studies by removing or modifying specific modules. Specifically, “w/o hetero models” excludes all additional heterogeneous base models from the stacking ensemble. “w/o HiLo GNNs” removes both high-pass and low-pass GNNs from the ensemble. “w/o Hi GNN” and “w/o Lo GNN” remove high-pass and low-pass GNN, respectively. “w/o hetero graph” removes all mixing transaction edges, retaining only address association edges in HAMIG. “w/o HiLo learning” replaces the high- and low-pass GNNs with two GCNs. “w/o label division” disables the confidence-based label division mechanism. “w/o stacking” discards the stacking framework and instead selects the better result between the high- and low-pass GNNs.
@@
As shown in Table 2, removing any component results in a consistent drop in both classification and ranking metrics, confirming the complementary contributions of all modules. Among all variants, excluding heterogeneous base models leads to the greatest degradation, with $\text{F}_{1}$ decreasing by 11.80% and MRR by 35.45%. Although the drop is substantial, this result is expected: in the w/o hetero models setting, the high-pass and low-pass GNNs are inherently homogeneous, limiting their ability to provide complementary information. In contrast, noticeable drops occur when the core GNNs are removed from the stacking ensemble, confirming their pivotal role in HiLoMix’s effectiveness. The remaining variants yield moderate performance decreases, further validating the necessity of each component. These smaller drops also demonstrate the robustness of HiLoMix and indicate that the stacking framework effectively integrates noise-resistant and complementary knowledge from heterogeneous base models, maintaining stable performance even when certain components are weakened.
@@
We conduct a comparative study across seven feature variants by progressively removing less important dimensions. Additionally, we incorporate node features from MixBroker for further comparison. As shown in Table 3, HiLoMix achieves the best performance across all metrics when using our 197-dimensional account node features, outperforming the second-best features, MixBroker features. Although the 221-dimensional feature set contains more attributes, its performance is inferior to that of the 197-dimensional set due to reduced interpretability. These results suggest that our 197-dimensional features not only enhance model explainability but also reduce HAMIG’s memory footprint without compromising performance.
@@
We perform a sensitivity analysis on key hyper-parameters in HiLoMix. We begin by evaluating model stability with respect to the number of folds used in cross-validation within the stacking ensemble. As shown in Figure 3, HiLoMix performance remains generally stable across different fold settings. The best overall performance is observed with 5-fold cross-validation, which is therefore adopted as the default setting. Both fewer and greater numbers of folds lead to slight performance degradation, potentially due to overfitting or unstable ensemble integration.
Figure 3: Performance of HiLoMix when training the base models in the stacking framework using different K of K-fold cross-validation.
@@
We further investigate two additional hyper-parameters: loss weighting $\lambda$, which balances the supervised and contrastive learning objectives, and the high-pass filter strength $\alpha$, which modulates the emphasis on high-frequency components. As illustrated in Figure 4, HiLoMix demonstrates robust performance across a wide range of $\lambda$ and $\alpha$ values. Both metrics peak around $\lambda=2$ and $\alpha=0.5$, suggesting that a moderate balance between the supervised and contrastive learning objectives and the well-calibrated filter strength are critical for optimal performance.
Efficiency Analysis
Although HiLoMix introduces additional components compared to simple baselines, most of them are computationally lightweight and do not incur substantial overhead. We compare HiLoMix with representative baselines, i.e., the best-performing model from each baseline category. The average training time per epoch is reported in Table 4.
Method Training time (s)
HiLoMix 6.91
GIN 0.26
NRGL 15.27
MixBroker 0.26
Table 4: Comparison of training efficiency.
As shown in Table 4, HiLoMix incurs moderate additional computation compared to simple GNNs but remains substantially faster than NRGL. In contrast, DeepWalk/Node2Vec and BERT4ETH require several hours for random walk generation or pretraining to obtain node embeddings, making them far less efficient than our model. All models in our experiments, including HiLoMix and the baselines, complete inference on the entire test set in under 0.01 seconds. In general, HiLoMix achieves a favorable balance between performance and computational cost.
Structural and Label Dynamics Analysis
[…]
Metric MIG HAMIG
Average node degree 0.0565 2.561
Average clustering coefficient 0.0004 0.0390
Graph density 4.93e-7 2.23e-5
Table 5: Comparison of graph structural properties between MIG and HAMIG.
We track the number of clean, confident-flipped, and remaining labels across epochs, as illustrated in Figure 5.
Figure 5: Dynamics of three label sets during training.
The evolution process can be roughly divided into two stages: early training (0–30 epochs) and later training (30–50 epochs). During the early stage, the number of clean labels gradually decreases, while remaining labels increase, as the model transitions from fitting clean samples to handling potentially noisy ones. At this point, confident-flipped labels remain nearly zero, indicating that the model has not yet learned sufficiently reliable structural patterns. In the later stage, the number of confident-flipped labels steadily increases, showing that the model begins to capture meaningful structural dependencies and distinguish true associations from incorrect ones. Meanwhile, clean labels stabilize, and the model progressively filters mislabeled samples from the remaining set, demonstrating improved robustness to noisy supervision. This dynamic evolution process confirms that confidence-based supervision effectively governs label noise and guides the model toward reliable learning.